Find F(x) for the following function.

f(x)=-4x+2

Then find F(5), f(0) and f(-7).

f(x)=___

f(5)=___

f(0)=___

f(-7)=___

Question complete.
Answer #1

Solution:

f(x) = -4x+2

f(5) = (-4*5)+2 = -20+2 = -18

f(0) = (-4*0)+2 = 0

f(-7) = (-4*-7)+2 = 28+2 = 30